OpenStudy (anonymous):

need help with # 57 and 61 i'll post the pic please and thank y

OpenStudy (anonymous):

you*

OpenStudy (unklerhaukus):

i dont really understand your question what did you get for #55 for example,

OpenStudy (unklerhaukus):

ah ok,

OpenStudy (unklerhaukus):

SO we want the tangent of t , in terms of sines only \[\tan(t)=\frac{\sin (t)}{\cos(t)}\] to get rid of the cosine , solve the trig identity \[\sin^2+\cos^2=1\]for cosine\[\quad\vdots\\\cos=\] substitute this result into \[\tan(t)=\frac{\sin (t)}{\cos(t)}\]

OpenStudy (anonymous):

the answer from the book is different :O

OpenStudy (anonymous):

sin t/ square root of 1-sin^2 t

OpenStudy (unklerhaukus):

wait , i think you have take into consideration that we are in the fourth quadrant

OpenStudy (anonymous):

okay

OpenStudy (unklerhaukus):

use this \[\begin{array}{r|l}S&A\\\hline T&C\end{array}\] sin,cos and tan are All positive in quadrant I only Sin is positive in quadrant II only tangent is positive in quadrant III only cosine is positive in quadrant IV

OpenStudy (anonymous):

its still confusing but thanks for trying to explain it to me :)

OpenStudy (unklerhaukus):

is sin t/ √( 1-sin^2 t) the answer you got or the answer in the book?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

both?

OpenStudy (anonymous):

that's the answer for # 57

OpenStudy (unklerhaukus):

yes, do you get it?

OpenStudy (anonymous):

nope

OpenStudy (unklerhaukus):

\[\sin^2+\cos^2=1\\ \cos^2=1-\sin^2\\\cos=\sqrt{1-\sin^2}\]

OpenStudy (anonymous):

i got that part

OpenStudy (unklerhaukus):

substitute this result into \[\tan(t)=\frac{\sin (t)}{\cos(t)}\]

OpenStudy (anonymous):

thats the part i dont get

OpenStudy (unklerhaukus):

so we have just showed that \[\cos(t)=\sqrt{1-\sin^2(t)}\] so \[\tan(t)=\frac{\sin (t)}{\cos(t)}=\frac{\sin(t)}{?}\]

OpenStudy (anonymous):

now i get it!!

OpenStudy (unklerhaukus):

good good

OpenStudy (unklerhaukus):

Do you think you can do 61?

OpenStudy (anonymous):

thanks